MAT 141 PLACEMENT TEST 9/13/02

NAME: ID number:

1. Let

f(x) = cos (x^{3}) g(x) = (cos x)^{3}.

Is f even, odd, or neither? Explain. Is g even, odd or neither? Explain.

2. Solve for x.

log_{5}x = log_{25}18 − log_{25}2.